OpenStudy (anonymous):

When 39,500 is divided by an integer n, the quotient is 123 and the remainder is 17. Find n.

OpenStudy (kainui):

Let's see if we can construct the formula for remainder and quotient and all that jazz by looking at some examples. Let's say 13 divided by 5 the quotient is 2 and the remainder is 3. So we can rewrite this as: 13/5=2+(3/5) See how this smaller version allows us to come up with a formula if we replace everything I called stuff like "quotient" is what we replace "2" but instead I'll use a "q" since it's shorter. x/n=q+r/n There we go. Try using that and coming up with it on your own if you have difficulty and I'll help.